# Mutual Information Optimization for SIM-Based Holographic MIMO Systems

### v1

@@ Abstract
In the context of emerging stacked intelligent metasurface (SIM)-based holographic MIMO (HMIMO) systems, a fundamental problem is to study the mutual information (MI) between transmitted and received signals to establish their capacity. However, direct optimization or analytical evaluation of the MI, particularly for discrete signaling, is often intractable. To address this challenge, we adopt the channel cutoff rate (CR) as an alternative optimization metric for the MI maximization. In this regard, we propose an alternating projected gradient method (APGM), which optimizes the CR of a SIM-based HMIMO system by adjusting signal precoding and the phase shifts across the transmit and receive SIMs in a layer-by-layer basis. Simulation results indicate that the proposed algorithm significantly enhances the CR, achieving substantial gains proportional to those observed for the corresponding MI. This justifies the effectiveness of using the channel CR for the MI optimization. Moreover, we demonstrate that the integration of digital precoding, even on a modest scale, has a significant impact on the ultimate performance of SIM-aided systems.
@@ I Introduction
Intelligent metasurfaces are expected to play a significant role in the evolution of future wireless communications. They are engineered surfaces, composed of a large number of controllable metamaterial elements that are capable of modifying the propagation of incident waves at the electromagnetic (EM) field level in dynamic and programmable manner [1]. This unique ability enables intelligent surfaces to improve energy efficiency, network capacity and connectivity, while also supporting other heterogeneous functions, such as localization and sensing. However, leveraging these advantages requires the acquisition of accurate channel state information (CSI), which presents a challenging task for conventional metasurface structures, such as reconfigurable intelligent surfaces. Furthermore, the multiplicative effect of the path loss of the RIS-assisted link severely reduces the possible gains through metasurface deployment. As a result, to maximize their potential, intelligent metasurfaces (e.g., RISs) are preferably placed in the vicinity of the transmitter and the receiver.
@@
The above discussions naturally motivate the integration of intelligent metasurface structures with wireless communication transceivers in massive MIMO (mMIMO) and HMIMO systems [2]. The use of intelligent metasurfaces in these systems offers a tradeoff between minimizing the number of radio frequency chains and maximizing the controllability over the radiated and received electromagnetic fields in the wave domain [3]. Moreover, the intelligent metasurfaces can provide beamforming capabilities comparable to those of conventional phased array antennas, but with much lower power consumption and cost. However, these metasurfaces are usually single-layer structures, which may limit their beamforming potential. In practical realizations, the elements of intelligent metasurfaces are constrained to a finite set of tunable states, and the number of these elements on a single metasurface is also limited. This can lead to beam misalignment, thereby undermining the expected performance results.
@@
To address the above limitations, multi-layer metasurface structures have emerged as a promising solution for flexibly forming different radiation patterns compared to their single-layer counterparts. Such structures, called SIMs, were recently introduced [4], which is inspired from the architecture of a deep neural network [5]. Indeed, SIMs mirrors the multi-layer structure of neural networks that are, if properly trained, capable of modeling various functions with improved signal processing capabilities. In the same way, when metasurface layers are properly placed, SIM can implement signal processing directly in the EM wave domain. This approach can potentially reduce the reliance on digital beamforming and high-precision digital-to-analog converters.
@@
In [6], the signal processing capabilities of SIMs were exploited for the implementation of 2D discrete Fourier transform (DFT) for direction of arrival (DOA) estimation. Moreover, a hybrid channel estimator, in which the received training symbols were processed first in the wave domain and later in the digital domain, was proposed in [7]. A general path loss model for an SIM-assisted wireless communication system was developed in [8], based on which, an algorithm for the receive power maximization was derived. In [9], the authors studied the achievable sum-rate maximization for a downlink channel between a SIM-assisted base station and multiple single-antenna users. Integrating SIMs with transmitters and receivers into the so-called SIM-based HMIMO system, which performs signal precoding and combining in the wave domain, was proposed in [10]. The introduced channel fitting approach enables the SIM-based HMIMO system to achieve significant channel capacity gains compared to mMIMO and RIS-assisted counterparts. Furthermore, the achievable rate optimization for the SIM-based HMIMO system was studied in [11].
@@
Despite quite a few papers about SIM-based systems, none of them has considered the achievable rate of such systems using discrete signaling, i.e., the MI. This gap has motivated us to investigate the optimization of the MI in the aforementioned systems. Since the direct optimization of the MI is intractable, we instead propose the use of the channel CR as an alternative metric to facilitate the MI optimization [12, 13]. The channel CR can be seen as a lower bound on the MI [12, Eq. (36)] and is widely recognized as a practical upper limit on the information rate that guarantees reliable communications.
@@
  • We show that using the channel CR as an alternative optimization metric enables efficient maximization of the MI in SIM-based HMIMO systems. Specifically, the CR is expressed as a closed-form expression, which facilitates the derivation of efficient optimization algorithms.
  • To maximize the CR, we formulate a joint optimization problem of transmit precoding and the phase shifts for both transmit and receive SIMs. Then an APGM is proposed to solve this problem, which particularly optimizes the phase shifts for the transmit and receive SIMs in a layer-by-layer basis. Moreover, we provide the computational complexity of the proposed algorithm.
@@ II System Model
We consider a SIM-based HMIMO communication system with $N_{t}$ transmit antennas and $N_{r}$ receive antennas, where both the transmitter and the receiver are equipped with SIMs. The transmit SIM consists of $L$ metasurface layers with $N$ meta-atoms in each layer, while the receive SIM consists of $K$ metasurface layers with $E$ meta-atoms in each layer. The meta-atoms of the SIMs are connected to smart controllers or field programmable gate array (FPGA) devices, which can independently adjust the phase shift of each meta-atom. In this way, MIMO transceivers with closely spaced SIMs can implement signal beamforming directly in the EM wave domain.
For the $l$-th layer of the transmit SIM, the phase shift matrix is denoted as $\boldsymbol{\mathbf{\Phi}}^{l}=\text{diag}(\boldsymbol{\phi}^{l})=\text{diag}([\phi_{1}^{l}\>\phi_{2}^{l}\>\cdots\>\phi_{N}^{l}]^{T})$, where $\text{diag}(\cdot)$ transforms a vector into a diagonal matrix. Also, $\phi_{n}^{l}=\exp(j\theta_{n}^{l})$, where $\theta_{n}^{l}$ is the phase shift of the $n$-th meta-atom in the $l$-th layer. Signal propagation between the $(l-1)$-th and $l$-th layer of the transmit SIM is modeled by the matrix $\mathbf{W}^{l}\in\mathbb{C}^{N\times N}$, where $[\mathbf{W}^{l}]_{m,n}$ denotes the signal propagation between the $n$-th meta-atom of the $(l-1)$-th layer and the $m$-th meta-atom of the $l$-th layer. According to Rayleigh-Sommerfeld diffraction theory, $[\mathbf{W}^{l}]_{m,n}$ is given by [14, Eq. (1)]
@@
for $l=2,3,\dots,L$, where $A$ is the area of each meta-atom, $d_{m,n}$ is the propagation distance between the meta-atoms in the $(l-1)$-th and $l$-th layer of the transmit SIM, $\chi_{m,n}$ is the angle between the propagation direction and normal direction of the $(l-1)$-th layer [^1] In this paper, it is assumed that all layers of the transmit SIM have the same arrangement and thus the right-hand side (RHS) of (1) does not depend on $l$. The same applies for the receive SIM.. Similarly, signal propagation between the transmit antenna array and the first layer of the transmit SIM is modeled by $\mathbf{W}^{1}\in\mathbb{C}^{N\times N_{t}}$, which is defined as (1). Finally, the EM wave domain beamforming matrix of the transmit SIM can be written as
$\mathbf{B}=\boldsymbol{\mathbf{\Phi}}^{L}\mathbf{W}^{L}\cdots\boldsymbol{\mathbf{\Phi}}^{2}\mathbf{W}^{2}\boldsymbol{\mathbf{\Phi}}^{1}\mathbf{W}^{1}\in\mathbb{C}^{N\times N_{t}}.$ (2)
At the receive SIM, the phase shift matrix for the $k$-th layer is given by $\boldsymbol{\mathbf{\Psi}}^{k}=\text{diag}(\text{$\boldsymbol{\psi}$}^{k})=\text{diag}([\psi_{1}^{k}\>\psi_{2}^{k}\>\cdots\>\psi_{E}^{k}]^{T})$, where $\psi_{e}^{k}=\exp(j\upsilon_{e}^{k})$ and $\upsilon_{e}^{k}$ is the phase shift of the $e$-th meta-atom in the $k$-th layer. Signal propagation between the $k$-th and $(k-1)$-th layer of the receive SIM is modeled by the matrix $\mathbf{U}^{k}\in\mathbb{C}^{E\times E}$, which is defined similarly as in (1). Also, signal propagation between the first layer of the receive SIM and the receive antenna array is given by $\mathbf{U}^{1}\in\mathbb{C}^{N_{r}\times E}$. The EM wave domain beamforming matrix for the receive SIM can be expressed as
@@
Let $\mathbf{G}$ be the channel between the transmit and the receive SIM. Then the end-to-end channel matrix for the considered communication system is written as
$\mathbf{H}=\mathbf{Z}\mathbf{G}\mathbf{B}\in\mathbb{C}^{N_{r}\times N_{t}}$ (4)
and the signal vector at the receive antennas is given by
@@ III Problem Formulation
where $\kappa_{i,j}=(-||\mathbf{H}\mathbf{P}(\mathbf{x}_{i}-\mathbf{x}_{j})+\mathbf{n}||^{2}+||\mathbf{n}||^{2})/\sigma^{2}$.
As the direct optimization of the MI in (7) involves discrete variables, which is intractable, we instead consider the optimization of a closed related metric, known as the CR, which is given by
@@
$\operatorname{subject~{}to}\text{ }$ $\bigl{|}\boldsymbol{\phi}\bigr{|}=1,$ (9b)
$\bigl{|}\boldsymbol{\psi}\bigr{|}=1,$ (9c)
@@ 39                                             $n\leftarrow n+1$
where $\nabla_{\mathbf{P}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ is the gradient of $f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ with respect to (w.r.t.) $\mathbf{P}^{*}$, $\mathcal{P}_{P}(\cdot)$ denotes the projection onto the set defined by (6) and $\nu_{n}$ is the step size, which is found by the line search routine described in steps 1 to 1 of 1.
In optimizing the phase shifts of the SIMs, existing algorithms typically perform a projected gradient step across all phase shifts of all layers simultaneously using the same step size [6, 7]. However, we have observed that such a method results in a poor performance. To address this, our proposed algorithm optimizes the phase shifts for the transmit and the receive SIMs in a layer-by-layer basis, with each layer being assigned a separate step size. This approach is shown to yield better performance results in our numerical experiments.
@@
$\nabla_{\mathbf{P}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ $=$ $-\frac{1}{4\sigma^{2}}\mathbf{H}^{H}\mathbf{H}\mathbf{P}\sum_{i,j=1}^{N_{vec}}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}$ (13)
$\nabla_{\boldsymbol{\phi}^{l}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ $=$ $-\frac{1}{4\sigma^{2}}\sum_{i,j=1}^{N_{vec}}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\operatorname{vec}_{d}(\mathbf{L}_{l})$ (14)
$\nabla_{\boldsymbol{\psi}^{k}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ $=$ $-\frac{1}{4\sigma^{2}}\sum_{i,j=1}^{N_{vec}}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\operatorname{vec}_{d}(\mathbf{K}_{k})$ (15)
@@
After calculating the gradients, the next step is to perform projection onto the corresponding feasible sets. It is easy to see from (6) that $\mathcal{P}_{P}(\mathbf{P})$ is given by
@@
$\overline{\phi_{n}^{l}}=\begin{cases}\phi_{n}^{l}/\bigl{|}\phi_{n}^{l}\bigr{|},&\phi_{n}^{l}\neq 0\\ \exp(j\alpha),\alpha\in[0,2\pi]&\phi_{n}^{l}=0.\end{cases}$ (18)
For $\phi_{n}^{l}=0$, the projection can take any point on the unit circle. Finally, $\mathcal{P}_{\boldsymbol{\psi}}(\boldsymbol{\psi}^{k})$ is obtained similarly as in (18).

### v2

@@ Abstract
In the context of emerging stacked intelligent metasurface (SIM)-based holographic MIMO (HMIMO) systems, a fundamental problem is to study the mutual information (MI) between transmitted and received signals to establish their capacity. However, direct optimization or analytical evaluation of the MI, particularly for discrete signaling, is often intractable. To address this challenge, we adopt the channel cutoff rate (CR) as an alternative optimization metric for the MI maximization. In this regard, we propose an alternating projected gradient method (APGM), which optimizes the CR of a SIM-based HMIMO system by adjusting signal precoding and the phase shifts across the transmit and receive SIMs on a layer-by-layer basis. Simulation results indicate that the proposed algorithm significantly enhances the CR, achieving substantial gains, compared to the case with random SIM phase shifts, that are proportional to those observed for the corresponding MI. This justifies the effectiveness of using the channel CR for the MI optimization. Moreover, we demonstrate that the integration of digital precoding, even on a modest scale, has a significant impact on the ultimate performance of SIM-aided systems.
@@ I Introduction
Intelligent metasurfaces are expected to significantly impact the evolution of future wireless communications. These surfaces, composed of a large number of controllable metamaterial elements, can dynamically modify the electromagnetic (EM) wave propagation [1]. This ability enables them to improve energy efficiency [2], network capacity, reliability [3] and connectivity [4], while also supporting other functionalities, such as localization and sensing. However, their advantages hinge on acquiring accurate channel state information (CSI), a challenging task for conventional structures, such as reconfigurable intelligent surfaces. Furthermore, the path loss in the RIS-assisted links can substantially diminish these benefits, indicating that placing intelligent metasurfaces close to transmitters and receivers maximizes their potential.
@@
The above discussions naturally motivate the integration of intelligent metasurface structures with wireless communication transceivers in massive MIMO (mMIMO) and HMIMO systems [5]. The use of intelligent metasurfaces in these systems offers a tradeoff between minimizing the number of RF chains and increasing control over EM fields [6]. Moreover, they can provide beamforming capabilities comparable to those of conventional phased array antennas, but at lower cost and power consumption. However, limitations due to their typically single-layer structure and the finite number of tunable states of their elements in practice can result in beam misalignment, potentially compromising the expected performance.
@@
To address the above limitations, multi-layer metasurface structures have emerged as a promising solution for flexibly forming different radiation patterns compared to their single-layer counterparts [7]. Such structures, called SIMs, were recently introduced [8], and their origin was inspired from the architecture of a deep neural network [9]. Indeed, SIMs mirror the multi-layer structure of neural networks that are capable of modeling various functions with improved signal processing capabilities. Similarly, when metasurface layers are properly placed, SIM can implement signal processing directly in the EM wave domain. This approach can potentially reduce the reliance on digital beamforming and high-precision digital-to-analog converters.
@@
In [10], the signal processing capabilities of SIMs were exploited for the implementation of 2D discrete Fourier transform (DFT) for direction of arrival (DOA) estimation. Moreover, a hybrid channel estimator, in which the received training symbols were processed first in the wave domain and later in the digital domain, was proposed in [11]. In [12], the authors studied the achievable sum-rate maximization for a downlink channel between a SIM-assisted base station and multiple single-antenna users. Integrating SIMs with transmitters and receivers into the so-called SIM-based HMIMO system, which performs signal precoding and combining in the wave domain, was proposed in [13]. Moreover, the achievable rate optimization for the SIM-based HMIMO system was studied in [14].
@@
Despite quite a few papers about SIM-based systems, none of them has considered the achievable rate of such systems using discrete signaling, i.e., the MI. This gap has motivated us to investigate the optimization of the MI in the aforementioned systems. As shall be seen shortly, the direct optimization of the MI is intractable, and thus we instead propose the use of the channel CR as an alternative metric to facilitate the MI optimization [15, 16], which is justified as follows. First, the MI is related to the CR, $R_{0}$, by $\text{MI}\geq N_{r}(1-\log_{2}e)+R_{0}$ [15, Eqn. (36)], meaning that maximizing the CR accordingly translates into maximizing the MI. Second, $R_{0}$ and the codeword error probability, $P_{e}$, are related by $P_{e}\leq 2^{-n(R_{0}-R)}$ [17, Eq. (6.8–19)], where $n$ is the codeword length and $R$ is the information rate. Essentially, letting $n\rightarrow\infty$, $P_{e}$ can be made arbitrarily small provided that $R<R_{0}$. Since the channel capacity represents the theoretical maximum information rate for reliable communications, the CR is effectively employed as a practical lower bound on the channel capacity. The inter-connection among the CR, the MI, and the channel capacity indeed suggests that the CR can be used to optimize the MI in an indirect, yet efficient, way.
@@
  • We show that using the channel CR as an alternative optimization metric enables efficient maximization of the MI in SIM-based HMIMO systems. Specifically, the CR is expressed as a closed-form expression, which facilitates the derivation of an efficient optimization algorithm.
  • To maximize the CR, we formulate a joint optimization problem of transmit precoding and the phase shifts for both the transmit and the receive SIMs. An APGM is proposed to solve this problem, which particularly optimizes the phase shifts for the transmit and the receive SIMs on a layer-by-layer basis, using an independent step size for each layer, which is different from [13]. Also, we provide the computational complexity of the proposed algorithm.
@@
Notation: Bold lower and upper case letters represent vectors and matrices, respectively. $\operatorname{Tr}(\cdot)$ denotes the trace of a matrix and $\text{diag}(\cdot)$ transforms a vector into a diagonal matrix. $\log_{2}(\cdot)$ is the binary logarithm, $\exp(\cdot)$ denotes the exponent, $\mathbb{E}\{\cdot\}$ denotes the expectation operation and $(\cdot)^{*}$ is the complex conjugate. $(\cdot)^{T}$ and $(\cdot)^{H}$ represent transpose and Hermitian transpose, receptively. $\nabla_{x}f(\cdot)$ is the complex gradient of $f(\cdot)$ with respect to (w.r.t.) $x^{*}$.
@@ II System Model
We consider a SIM-based HMIMO communication system with $N_{t}$ transmit antennas and $N_{r}$ receive antennas, where both the transmitter and the receiver are equipped with SIMs. The transmit SIM consists of $L$ metasurface layers with $N$ meta-atoms in each layer, while the receive SIM consists of $K$ metasurface layers with $E$ meta-atoms in each layer. We assume that the signal propagation through every SIM layer is ideal, i.e., without any power loss. The SIM meta-atoms are connected to smart controllers, which can independently adjust the phase shift of each meta-atom. In this way, MIMO transceivers with closely spaced SIMs can implement signal beamforming directly in the EM wave domain.
For the $l$-th layer of the transmit SIM, the phase shift matrix is denoted as $\boldsymbol{\mathbf{\Phi}}^{l}=\text{diag}(\boldsymbol{\phi}^{l})=\text{diag}([\phi_{1}^{l}\>\phi_{2}^{l}\>\cdots\>\phi_{N}^{l}]^{T})$. Also, $\phi_{n}^{l}=\exp(j\theta_{n}^{l})$, where $\theta_{n}^{l}$ is the phase shift of the $n$-th meta-atom in the $l$-th layer. Signal propagation between the $(l-1)$-th and $l$-th layer of the transmit SIM is modeled by the matrix $\mathbf{W}^{l}\in\mathbb{C}^{N\times N}$, where $[\mathbf{W}^{l}]_{m,n}$ denotes the signal propagation between the $n$-th meta-atom of the $(l-1)$-th layer and the $m$-th meta-atom of the $l$-th layer. According to Rayleigh-Sommerfeld diffraction theory, $[\mathbf{W}^{l}]_{m,n}$ is equal [18, Eq. (1)]
@@
for $l=2,3,\dots,L$, where $A$ is the area of each meta-atom, $\lambda$ is the wavelength, $d_{m,n}$ is the propagation distance between the meta-atoms in the $(l-1)$-th and $l$-th layer of the transmit SIM, $\chi_{m,n}$ is the angle between the propagation direction and normal direction of the $(l-1)$-th layer [^1] In this paper, it is assumed that all layers of the transmit SIM have the same arrangement and thus the right-hand side (RHS) of (1) does not depend on $l$. The same applies to the receive SIM.. Also, signal propagation between the transmit antenna array and the first layer of the transmit SIM is modeled by $\mathbf{W}^{1}\in\mathbb{C}^{N\times N_{t}}$, which is defined as (1). The wave domain beamforming matrix of the transmit SIM is given as $\mathbf{B}=\boldsymbol{\mathbf{\Phi}}^{L}\mathbf{W}^{L}\cdots\boldsymbol{\mathbf{\Phi}}^{2}\mathbf{W}^{2}\boldsymbol{\mathbf{\Phi}}^{1}\mathbf{W}^{1}\in\mathbb{C}^{N\times N_{t}}.$
@@
Let $\mathbf{G}\in\mathbb{C}^{N\times E}$ be the channel between the transmit and the receive SIM. Then the end-to-end channel matrix for the considered communication system is $\mathbf{H}=\mathbf{Z}\mathbf{G}\mathbf{B}\in\mathbb{C}^{N_{r}\times N_{t}}$ and the signal vector at the receive antennas is given by
@@ III Problem Formulation
where $\kappa_{i,j}=(-||\mathbf{H}\mathbf{P}(\mathbf{x}_{i}-\mathbf{x}_{j})+\mathbf{n}||^{2}+||\mathbf{n}||^{2})/\sigma^{2}$. As the direct optimization of the MI in (4) involves discrete variables, which is intractable, we instead consider the optimization of a closed related metric, known as the CR, which is given by
@@
$\operatorname{subject~{}to}\text{ }$ $\bigl{|}\phi_{n}^{l}\bigr{|}=1,\;\forall l,n$ (6b)
$\bigl{|}\psi_{e}^{k}\bigr{|}=1,\;\forall k,e$ (6c)
@@ 39                                             $n\leftarrow n+1$
where $\nabla_{\mathbf{P}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ is the gradient of $f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ w.r.t. $\mathbf{P}^{*}$, $\mathcal{P}_{P}(\cdot)$ denotes the projection onto the set defined by (3) and $\nu_{n}$ is the step size, which is found by the line search routine described in steps 1 to 1 of 1.
In optimizing the phase shifts of the SIMs, existing algorithms typically perform a projected gradient step across all phase shifts of all layers simultaneously using the same step size [10, 11]. However, we have observed that such a method results in poor performance (see Fig. 1). To address this, our proposed algorithm optimizes the phase shifts for the transmit and the receive SIMs on a layer-by-layer basis, with each layer being assigned a separate step size. This approach is shown to yield better performance results in our numerical experiments.
@@
$\nabla_{\mathbf{P}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ $=-\frac{1}{4\sigma^{2}}\mathbf{H}^{H}\mathbf{H}\mathbf{P}\sum\nolimits_{i,j=1}^{N_{vec}}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}$ (10)
$\nabla_{\boldsymbol{\phi}^{l}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ $=-\frac{1}{4\sigma^{2}}\sum\nolimits_{i,j=1}^{N_{vec}}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\operatorname{vec}_{d}(\mathbf{L}_{l})$ (11)
$\nabla_{\boldsymbol{\psi}^{k}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ $=-\frac{1}{4\sigma^{2}}\sum\nolimits_{i,j=1}^{N_{vec}}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\operatorname{vec}_{d}(\mathbf{K}_{k})$ (12)
@@
After calculating the gradients, the next step is to perform a projection onto the corresponding feasible sets. It is easy to see from (3) that $\mathcal{P}_{P}(\mathbf{P})$ is given by
@@
$\overline{\phi_{n}^{l}}=\begin{cases}\phi_{n}^{l}/\bigl{|}\phi_{n}^{l}\bigr{|},&\phi_{n}^{l}\neq 0\\ e^{j\alpha},\alpha\in[0,2\pi]&\phi_{n}^{l}=0.\end{cases}$ (15)
For $\phi_{n}^{l}=0$, the projection can take any point on the unit circle. Finally, $\mathcal{P}_{\boldsymbol{\psi}}(\boldsymbol{\psi}^{k})$ is obtained similarly as in (15). The convergence proof of the proposed method follows the same arguments used in [19] and is omitted for brevity.